A manufacturing business is considering investing in some new equipment. The management accountant has estimated the future net cash flows from the investment as follows.

Initial investment

(£1,360,000)

Year 1

£470,000

Year 2

£580,000

Year 3

£580,000

Year 4

£500,000

This business uses straight-line depreciation and its cost of capital (the discount rate for investment appraisal is 10%). It is assumed that the new equipment will have a residual value of zero at the end of four years.

Required:

  1. Calculate the payback period,
  1. accounting rate of return (ARR) and

(6 marks)

(6 marks)

  1. the net present value (NPV) for the proposed investment.

(6 marks)

  1. Advise the business whether, on financial grounds, it should invest in the new equipment.

(2 marks)

  1. Explain why the NPV method is considered superior to the payback period and accounting rate of return methods of appraising potential capital investments.

(10 marks)

  1. Explain the effect of non-financial/non-quantitative factors on project appraisal by using relevant examples.

(10 marks) (Question 3 total: 40 marks)
Answer #1
a) Payback Period
Year Cash Flow Accumulated Cash Flow
£                        -   -£         13,60,000 -£                          13,60,000
£                          1 £            4,70,000 -£                            8,90,000
£                          2 £            5,80,000 -£                            3,10,000
£                          3 £            5,80,000 £                             2,70,000
£                          4 £            5,00,000 £                             7,70,000
payback petiod = 2 years + 310000/580000
=2.53years
b) Annual Depreciation = Intial investment - salavage value / life of the asset
=(1360000-0)/4
=340000 per year
Calculation of annual net income
Net Income = Annual cash flow - annual depreciation
Year Cash Flow Depreciation Net Income
£                    1.00 £      4,70,000.00 £                       3,40,000.00 £         1,30,000.00
£                    2.00 £      5,80,000.00 £                       3,40,000.00 £         2,40,000.00
£                    3.00 £      5,80,000.00 £                       3,40,000.00 £         2,40,000.00
£                    4.00 £      5,00,000.00 £                       3,40,000.00 £         1,60,000.00
Total £         7,70,000.00
Average Net Income = 770000/4 years
=192500
Average Assets = (Opeaing Assets+Salvage Value) /2
=(1360000-0)/2
=680000
Accounting rate of return= average net income / average investment
=192500/680000
=28.31%
c) Net Present value
Year Cash Flow PV factor PV of cash flow
0 -£   13,60,000.00 £                                     1.00 -£      13,60,000.00
1 £      4,70,000.00 £                                     0.91 £         4,27,272.73
2 £      5,80,000.00 £                                     0.83 £         4,79,338.84
3 £      5,80,000.00 £                                     0.75 £         4,35,762.58
4 £      5,00,000.00 £                                     0.68 £         3,41,506.73
NPV £         3,23,880.88
d) Since the NPV of the new equipment is positive, the company should invest in the equipment.